If one root of the equation x^2+px+q=0 is twice of the second root, then prove that 2p^2=9q.
Answer by Boreal(15235)

quadratic formula
x=(1/2)(-p+sqrt(p^2-4q)) and that is 2*(1/2) (-p-sqrt (p^2-4q))
multiply both sides by 2
-p+sqrt(p^2-4q)=2*(-p-sqrt(p^2-4q))
-p+sqrt(p^2-4q)=-2p-2sqrt(p^2-4q)
p=-3 sqrt(p^2-4q)
square both sides
p^2=9 (p^2-4q)=9p^2-36q
-8p^2= -36 q
2p^2=9q